OpenStudy (anonymous):

Using the point-slope formula, what is the equation of the line that passes through the points (1,5) and (0,0)?

OpenStudy (mertsj):

Can you find the slope?

OpenStudy (anonymous):

i just need to know the point-slope formula then i think i can do this

OpenStudy (mertsj):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (anonymous):

so it is 5-1 = m(1-0)

OpenStudy (mertsj):

If you are writing the equation of a line you need the variables x and y so keep them and only substitute for the x1 and y1

OpenStudy (anonymous):

so y=5x-1

OpenStudy (mertsj):